OpenStudy (anonymous):

if a number is added to twice the number, the result is -12. what is the number

OpenStudy (anonymous):

x + 2x=-12 3x=-12 x=-4

OpenStudy (saifoo.khan):

x + 2x = -12 Solve.

OpenStudy (anonymous):

x+2x= -12 3x =-12 x= -12/3 x=-4